5. Calculate the energy necessary to excite an electron from the second shell to fifth shell in a hydrogen atom, in kJ/mol.
Answer #1

= 13.6 z [ tai ev = 13.6** [*3] ev = 13-6 * [5 - 3s]. er = 13.6 * 25-4 ev 25X4 = 13.6 x 21 ev 1oo = 2.856 ev ser = 1.602 x 1